OpenStudy (anonymous):

What is an equation of the line in slope-intercept form? m = two-sevenths and the y-intercept is (0, –12)

ganeshie8 (ganeshie8):

u knw wat a slope intercept form for a line is ?

OpenStudy (anonymous):

isn't it that line that curves?

ganeshie8 (ganeshie8):

slope intercept form, is just an equation, as the question says

ganeshie8 (ganeshie8):

here is the slope intercept form :- y = mx + b

ganeshie8 (ganeshie8):

m = slope b = y-intercept

ganeshie8 (ganeshie8):

just plug the given info into that :)

OpenStudy (anonymous):

so I would jst flip it around?

ganeshie8 (ganeshie8):

m = two-sevenths and the y-intercept is (0, –12) m = 2/7, b = -12

OpenStudy (anonymous):

so I just take out the 0?

ganeshie8 (ganeshie8):

slope intercept form :- y = mx +b plug the values, y = 2/7 x - 12

ganeshie8 (ganeshie8):

we're done. i was not following wat u saying sorry lol -_-

ganeshie8 (ganeshie8):

thought, if i showed u the complete working, u wud understand the glaringly obvious thing ;)

OpenStudy (anonymous):

Thats ok. I have another question like this one..but it is (8, –8); slope: 3 and another one y = –1.7x + 8.5 would I do the same to these two as well?

OpenStudy (anonymous):

Yes, and it helped thank you for showing me

ganeshie8 (ganeshie8):

write the complete question

ganeshie8 (ganeshie8):

it is (8, –8); slope: 3

OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = –1.7x + 8.5 What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3

ganeshie8 (ganeshie8):

ohkay, i forgot how a standard form looks like for a line, can u google and get me, the "Standard Form of Line"

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

ty il wait :)

OpenStudy (anonymous):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

ganeshie8 (ganeshie8):

Ahh yes... standard form is :- Ax + By = C

ganeshie8 (ganeshie8):

given equation :- y = –1.7x + 8.5

ganeshie8 (ganeshie8):

what should be our first step to change it to Standard form.. any guess ? :)

OpenStudy (anonymous):

add -1.7 to 8.5?

ganeshie8 (ganeshie8):

yup, add -1.7x to both sides. DONT IGNORE THE X

OpenStudy (anonymous):

1.7y and 10.2

ganeshie8 (ganeshie8):

y = –1.7x + 8.5 add 1.7x both sides y + 1.7x = 8.5 1.7x + y = 8.5

ganeshie8 (ganeshie8):

you cant add apples and oranges

ganeshie8 (ganeshie8):

you cant add 1.7x to 8.5

ganeshie8 (ganeshie8):

below complete working for converting it to standard form :- y = –1.7x + 8.5 add 1.7x both sides y + 1.7x = 8.5 1.7x + y = 8.5 multiply 10 thru out 17x + 10y = 85

ganeshie8 (ganeshie8):

Now does it look like Ax + By = C ?

OpenStudy (anonymous):

yes and it looks like one of the aswers too

ganeshie8 (ganeshie8):

cool /

OpenStudy (anonymous):

thank you for the help, once again :)

ganeshie8 (ganeshie8):

np :) u able to do the other problem ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good :)